OpenStudy (anonymous):

A landscaping company has two orders for bushes and trees. Order 1 for 12 bushes and 15 trees that will cost the customer $600 Order 2 for 8 bushes and 20 trees that will cost the customer $680 What is the cost of one bush?

OpenStudy (mathstudent55):

First, choose a variable to represent the number of bushes, and another variable to represent the number of trees.

OpenStudy (aum):

Let the cost of 1 bush be x and the cost of 1 tree be y. "12 bushes and 15 trees that will cost the customer $600" 12x + 15y = 600 ----- (1) "8 bushes and 20 trees that will cost the customer $680" 8x + 20y = 680 ------ (2) Eliminate y from (1) and (2) and solve for x.

OpenStudy (anonymous):

Thank you.

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

Wait. so the answer is..?

OpenStudy (aum):

12x + 15y = 600 ----- (1) This can be simplified by dividing by 3: 4x + 5y = 200 8x + 20y = 680 ------ (2) This can be simplified by dividing by 4: 2x + 5y = 170 So we have: 4x + 5y = 200 2x + 5y = 170 To eliminate y, subtract the two equations. What do you get?

OpenStudy (aum):

If you subtract the two equations, the 5y will cancel and you get 2x = 30 x = 15 So the price of one bush = $15.

OpenStudy (anonymous):

Thank you so much.

OpenStudy (aum):

You are welcome.